

********************************************************************
UARS WEEKLY STATUS REPORT - 07 Nov 1997
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems. A lamp
calibration scan was also successfully performed.
***********
PEM
***********
PEM zenith spectrometers were active from October 31 through November 2.
During this time, PEM produced energy deposition profiles of the electron and
proton precipitation.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
HALOE remains in survival mode. HALOE is scheduled to be powered on
beginning-of-day on November 11, 1997.
***********
MLS
***********
MLS was reactivated as scheduled on Oct 31 @18:50, in the reduced 85 watt power
mode. MLS was then powered off at 23:43 due to low EON voltage. MLS was
reactivated in the 85 watt power mode again on Nov 2 @02:52. MLS was again
powered off due to low EON voltage on Nov 2 @09:54. MLS was reactivated in the
85 watt power mode on Nov 2 @12:20 and made full vertical scans until MLS was
powered off on Nov 4 at 20:58 UT. MLS is scheduled to start north viewing
operations on Nov 20.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
UARS CDHF Status Week Ending November 7, 1997
---------------------------------------------
The software that generates the UARS level 0 files will be re-hosted to
a more current hardware platform on Saturday November 15, 1997. UARS CDHF users
should notice no change in the level 0 product. However, if users notice
anything unusual in the level 0 product or associated RAC transfer activities
please contact the PAC office immediately. A NEWS message will be released on
the CDHF with more detail about this activity.
The CDHF has received a total of 340 CD-ROM requests as of 07 November-1997.
One new CD-ROM order was received and shipped this week.
The CDHF successfully tested MLS V4.9 Engineering Software. Additional days
in Engineering Mode will be processed before MLS makes a final delivery and
production mode begins.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
30 days PEM (3.32) Prod (1350R-540C=810L)
27 days MLS (4.9) ENG (40R-27C=13L)
Pending CDHF software upgrades include IDL 5.0, and OpenIngres 2.0
***********
FOT
***********
REPORTING PERIOD
This report covers 24 October 1997 (Orbit 33443, GMT Day 298) through
31 October 1997 (Orbit 33548, GMT Day 304).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 298/2311 33458 Off
HRDI 99/0048 33459 On
ACRIM 300/0052 33474 On
SUSIM 300/1329 33482 On
HALOE 303/2326 33533 Off
HRDI 303/2326 33533 Standby
PEM 304/0147 33534 ZEPS On
MLS 304/1849 33545 On
MLS 304/2340 33548 Off
The beta angle is now decreasing toward a minimum angle of 0 degrees on
07 November 1997 (DOY 311).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery
1 remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.5 end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery
2 and Battery 3 of 1.3 degrees C. Additional battery performance data
for each battery on 25 October 1997 and 31 October 1997 are provided in
Appendix D.
The clock error ranged from -10.7 to +10.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 16
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% ACRIM 01
% HALOE 01
% HRDI 05
% MLS 03
% PEM 01
% SOLSTICE 01
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 04
TRANSPONDER FREQUENCIES:
Transponder A: 97/287 - 2287.496931
Transponder B: 97/295 - 2287.499126
GROUND SYSTEM OPERATION
Four (4) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 97/303 show the observatory data loss to
be 20 hours, 41 minutes, 03.985 seconds (an increase of 0.004 seconds).
This is a 0.0386 percent data loss which equals a 99.9614 percent data
capture for the mission.
The increase of 0.004 seconds occurred on Day 298 and was the result of
hits in the data.
The FOT observed that the SSPP was apparently not correctly tracking
star number 48000001 during several real-time events on 29 October 1997.
The SSPP beta angle in the real-time telemetry was approximately 13
degrees offset from the predicted value in the CMS generated tracking
reports. Investigation revealed an erroneous right ascension precession
rate in the latest FDF supplied star catalog. This error resulted in
the CMS software generating incorrect star position coordinates in the
weekly OBC System Table 65 load. Therefore, when the SSPP was commanded
(via stored commands) to track star 48000001, it was pointing at an
incorrect right ascension as defined by System Table 65. FDF was
notified and is currently working on generation of a new star catalog
with a correct precession rate for star 48000001. SOLSTICE operations
has been asked not to request SSPP to track the star in question.
The functionality of all remaining operational functions on UMPG Alpha
computer (node UMPGAL) was verified. Several DCL command file changes
were required to support printing of FOT plot files from UMPGAL. UMPGAL
will replace UMPG as the operational MPG node (for FOT purposes only)
late next week. The MPG plans on switching Daily Activity Plan (DAP)
processing to UMPGAL before the end of the November 1997.
No testing of the new CMS software release (R7B1V2) occurred this week.
The FOT and UCMS system managers are currently trying to find a spare
workstation which can be hooked up to the UCMSB (back-up) system. This
workstation is necessary for the FOT to test the next CMS software
release (R7B1V2), which currently resides on UCMSB in a test
environment.
FOT continues to assess participation in the IP return-link telemetry
test that the network is conducting through the end of November 1997.
Since the FOT is currently in the process of transition back to the
remodeled control center, UARSU involvement in the testing will most
likely be delayed until after 10 November 1997.
Several POCC database modifications were made and a preliminary version
of UARS Database Version 64 was delivered to the MSOCC software
coordinator this week. The modifications were necessary to provide more
accurate monitoring of the UARS yaw-around maneuvers. The new database
will be placed on the UTTS AP disk pack and tested by the FOT beginning
03 November 1997. Assuming no problems are found, the database will be
officially released next week and used for the upcoming yaw-around
maneuver on 07 November 1997.
Investigation of the stored command verification problem - time to
complete - associated with Applications Processor (AP) software release
10.08 continued. The cause has been identified, but rectifying the
problem continues.
OTHER SIGNIFICANT EVENTS
Monitoring of the magnetic field model performance continues.
FOT continues to develop an updated Power Profile for UARS due to the
current status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering
effort. The FOT delivered an ASCII format version of the OBCSTF
resident POCC database to the TOMS/UARS re-engineering team this week.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not
been scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 07 November
1997 at 311/132708.
The UARS orbit altitude is dropping slowly and a drag makeup orbit
adjust will probably be required in mid-1998.
FACILITIES
CRTs, computers, and other equipment are being installed in the semi-
circular console within the UARS/TOMS Mission Operations Room (MOR).
Operations continue from a dedicated room in the EOS area which is
across the hall.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 31 October 1997
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
33452 298/1258 97-529 HRDI Increase HTR delay - 600 Amp Min
33458 298/2311 97-530 MLS Power OFF
33459 299/0048 97-531 HRDI Power ON
33459 299/0048 97-532 HRDI Set Htr Delay Cycle - 400 amp min
33474 300/0052 97-533 ACRIM Power ON
33474 300/0052 97-534 SOLSTICE Microloads
33482 300/1329 97-535 SUSIM Power ON
33483 300/1516 97-536 OBC Clk Adj fr 20.6 to 0.00 msec/day
33486 300/2015 97-537 PSU Reset CH18 HTR cycle to Normal
Configuration
33512 302/1345 97-538 OBC/CLK Clk Adj fr 0.00 to 20.6 msec/day
33533 303/2326 97-539 HALOE Power OFF
33533 303/2326 97-540 HRDI HRDI to Standby
33534 304/0147 97-541 PEM "Power ON, ZEPS"
33542 304/1350 97-542 OBC/CLK Clk Adj fr 20.6 to 0.00 msec/day
33545 304/1849 97-543 MLS Power ON
33547 304/2154 97-544 HRDI HRDI to Standby again
33548 304/2340 97-545 MLS Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 31 October 1997
The following Out-of-Limit (ALARM) conditions were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
OBSVLBEON Y-LO Load Bus Volts at EON 33444-52,85,8,9,518
R-LO 33451,4,5,7,519-48
OBSBATVEON2 Y-LO Battery 2 voltage at EON 33444-52,5,89-518
R-LO 33451,4,7,519-48
OBSBATVEON3 Y-LO Battery 3 voltage at EON 33444-52,5,89-518
R-LO 33451,4,5,7,519-48
OBSSOC2 Y-LO Battery 2 Current SOC 33451,66,81,95-7,542,7,8
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 33444-49,51-84
QTSSPPPX Y-HI SSPP +X Thermostat 33451,66
QTSSPPMX Y-HI SSPP -X Thermostat 33480
HRIFESURVTMP Y-HI HRDI IFE Survival Temp 33535
HRRADTEMP Y-HI HRDI radiator temp 33535
The following Out-of-Limit (ALARM) conditions were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
ACTEMPANA1 Y-LO Temp PCU Board 33474-81,83-85,87-90
ACTEMPANA2 Y-LO Temp Sensor Electronics 33474-81,83-86
ACTEMPHM Y-LO Temp Heatsink Mass 33474-81
ACTEMPSA1 Y-LO Temp ACR 1 33474-77
ACTEMPSA2 Y-LO Temp ACR 2 33474-76
ACTEMPSA3 Y-LO Temp ACR 3 33474-77
ACTEMPSS Y-LO Temp Sun Sensor 33474-81
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 31 October 1997
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
33484 300/15:29 (27 Oct 97) from 20.6 to 00.0 msec/day
33512 302/13:50 (29 Oct 97) from 00.0 to 20.6 msec/day
33542 304/13:54 (31 Oct 97) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 31 October 1997
25 Oct. 1997 - GMT Day 298 Beta = 42.6 deg, SA Pos.= 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.68 +0.0/ +0.0 18.7 0.000 0.0
2 +89.6/ +0.0 +7.09/+5.79 +29.6/-16.8 24.5 1.020 23.5
3 +67.2/ -11.2 +5.47/+4.51 +28.8/-16.4 24.5 1.023 22.8
31 Oct. 1997 - GMT Day 304 Beta = 26.3 deg, SA Pos.= 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.05 +0.0/ +0.0 18.7 0.000 0.0
2 +78.4/ +0.0 +8.07/+7.09 +32.8/-16.4 23.7 1.017 23.9
3 +67.2/ -5.6 +7.09/+5.47 +31.6/-16.4 23.5 1.019 23.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 31 October 1997
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-111 FOT 33460 299/0228 14170 Generic Late
Acquisition
97-111 FOT 33472 299/2152 14170 Generic Late
Acquisition
97-118 FOT 33518 302/2322 19566 No acq. for entire
event
97-111 FOT 33516 302/2024 14171 Generic Late
Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-112 FOT 33110 275/1654 AP DOWN
97-113 FOT 33139 277/1508 YAW SKEW WHEEL DIR
97-114 FOT 33210 282/0944 19451 NO ODM OR GCMRUS
97-115 FOT 33262 285/2028 AP CRASHED
97-116 FPT 33263 285/2239 TAC CRASHED
97-117 FOT 33360 292/0935 19551 TDRS SSA FAILURE
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period